OpenStudy (anonymous):

I need help with trigonometric identities, I'm really confused with this questions. Derive this identity from the sum and difference formulas for cosine: sin a sin b = (1 / 2)[cos(a – b) – cos(a + b)] I need the claculation and reason.

ganeshie8 (ganeshie8):

Question says : `Derive this identity from the sum and difference formulas for cosine`

ganeshie8 (ganeshie8):

do u remember the sum/difference formulas for cosine ?

OpenStudy (anonymous):

Do u know that cos(a - b) = cos(a)cos(b) + sin(a)sin(b) cos(a + b) = cos(a)cos(b) - sin(a)sin(b)

OpenStudy (anonymous):

@ShantelT1

OpenStudy (anonymous):

No sorry I don't really understand this.

OpenStudy (anonymous):

These are the sum/difference formulas for cosine. The first formula is the sum and the second one is difference. You just learn them for now.

OpenStudy (anonymous):

Now u have to just subtract the second equation from the first,and u will get what u need....